Need help solving 3v + 2/3=v/6
Answer by stanbon(75887)

(3v + 2)/3 = v/6
----------
Cross-multiply to get:
6(3v+2) = 3v
18v + 12 = 3v
15v = -12
v = -4/5
================
Cheers,
Stan H.
================
-----